# Exploring the parameter dependence of atomic minima with implicit differentiation

### v1

@@ Abstract
Interatomic potentials are essential to go beyond ab initio size limitations, but simulation results depend sensitively on potential parameters. Forward propagation of parameter variation is key for uncertainty quantification, whilst backpropagation has found application for emerging inverse problems such as fine-tuning or targeted design. Here, the implicit derivative of functions defined as a fixed point is used to Taylor expand the energy and structure of atomic minima in potential parameters, evaluating terms via automatic differentiation, dense linear algebra or a novel sparse operator approach. The latter allows efficient forward and backpropagation through relaxed structures of arbitrarily large systems. The implicit expansion accurately predicts lattice distortion and defect formation energies and volumes with classical and machine-learning potentials, enabling high-dimensional uncertainty propagation without prohibitive overhead. We then show how the implicit derivative can be used to solve challenging inverse problems, minimizing an implicit loss to fine-tune potentials and stabilize solute-induced structural rearrangements at dislocations in tungsten.
Figure 1: Schematic of the implicit derivative approach in atomic systems. a A local minima ${\bf X}^{*}$ will change under changes in potential parameters ${{\bm{\Theta}}}$. The implicit derivative is the gradient of this change, allowing a first-order prediction of structural variations and a second-order prediction of energy variations. b Uncertainty quantification application. Having the uncertainty in the parameter space, $\Delta{{\bm{\Theta}}}$, implicit derivative allows one to compute the uncertainty of any molecular statics property $\mathcal{Q}$. c Inverse design application. With initial configuration ${\bf X}^{*}_{{{\bm{\Theta}}}_{1}}$ and potential ${{\bm{\Theta}}}_{1}$, one defines a target configuration ${\bf X}^{*}_{{{\bm{\Theta}}}_{2}}$. Implicit derivative allows one to effectively minimize the potential parameters that provide the target configuration.
@@ I Introduction
Regardless of the interatomic potential employed, changing parameters will change any quantity of interest extracted from a simulation. For simple classical models, parameter variation is essential to explore the model’s phenomenology [7, 5]. For modern data-driven models which target quantitative accuracy, parameter uncertainties can be estimated by Bayesian regression on training data [8, 9] and should be forward propagated to simulation results to bound quantities of interest [10].
@@
Any forward propagation scheme must account for the strong correlation between individual energy or force evaluations when calculating e.g. formation energies or dynamical averages, as these will strongly affect (typically reduce) uncertainty on the final simulation result. Backpropagation of structural modifications to changes in parameters is finding application in training interatomic potentials from experimental data [11] or tuning simple interatomic potentials to reproduce desired self-assembly kinetics [7]. More recently, ‘universal’ machine learning potentials have shown near-quantitative accuracy across large portions of the periodic table [12, 13]. This has raised interest in back-propagation for fine-tuning universal models for specific applications [14], and opened the possibility of targeted design through navigation of a smooth latent composition representation [15].
@@
Forward and backpropagation of parameter variation through complex simulations is typically achieved using reverse-mode automatic differentiation (AD) routines, which offer high arithmetic efficiency at the price of large memory requirements  [16, 17, 18]. Whilst AD offers clear advantages in implementation, their significant memory burden complicates application to large atomic systems, especially when targeting higher derivatives beyond forces. Existing methods to propagate variation in parameters to variation in simulation results thus typically employ resampling: new parameters are drawn from some distribution [10] and simulations are repeated. Whilst conceptually straightforward, back-propagation is not possible, assessing convergence is challenging, and the cost can be potentially very large, especially for simulation results requiring geometry minimization for each potential sample.
@@
In this paper, we explore an alternative approach, analytically expanding the structure of relaxed minima to first order in parameter variation, giving a second order expansion of the total energy. The expansion is achieved through evaluation of an implicit derivative [19], i.e. the derivative of a function defined as fixed point, in this case, the atomic structure of a local minimum (figure 1a). Our main results are that the implicit derivative enables 1) forward propagation of parameter uncertainties to simulation results for orders of magnitude less computational effort than resampling schemes, allowing rapid propagation of parameter uncertainties and 2) back-propagation of structural variations to target composition-induced structural rearrangements in multi-thousand-atom systems, a challenging task for any other approach (see illustration of these two ideas in figures 1b and 1c).
@@
We implement and compare methods to evaluate the implicit derivative using AD [17, 18], dense linear algebra and a novel sparse operator approach, using the jax-md [16] and LAMMPS [20] simulation codes. We find AD routines for the implicit derivative reach GPU memory limits for $\sim$1000-atom systems even on best-in-class A100 hardware. In contrast, our sparse operator reduces to a constrained minimization in LAMMPS, allowing memory-efficient and highly parallelized evaluation. This innovation allows uncertainty quantification and inverse design studies with the large atomic simulations essential to capture realistic defect structures.
@@
The paper is structured as follows. We first define the implicit derivative, the Taylor expansion approximations used and their evaluation using automatic differentiation or linear algebra techniques. We then describe how the implicit derivative can be used in forward and backpropagation of parameter variations. Forward propagation of parameter variation is demonstrated using classical and machine learning potentials to explore lattice distortion and vacancy defect formation [28]. Finally, we demonstrate back-propagation of parameter variations, finding parameter variations which stabilize subtle solute-induced dislocation core reconstructions in tungsten [4].
@@ II.1 Implicit derivative of atomic configurations
We consider a system of $N$ atoms in a periodic supercell of volume $V$, with atomic coordinates ${\bf X}\in\mathbb{R}^{N\times 3}$ and a supercell matrix ${\bf C}\in\mathbb{R}^{3\times 3}$, $V={\rm Det}({\bf C})$. Changes to the supercell ${\bf C}\to{\bf C}+\delta{\bf C}$ are defined to induce homogeneous deformations, as in e.g. energy volume curves. As a result, we work with scaled atomic coordinates $\tilde{\bf X}$, such that ${\bf X}\equiv\tilde{\bf X}{\bf C}$ and the tuple $(\tilde{\bf X},{\bf C})$ fully defines atomic configurations with periodic boundary conditions and fixed atomic species. With a vector of $N_{D}$ potential parameters ${{\bm{\Theta}}}$, a potential energy model ${U}(\tilde{\bf X},{\bf C};{{\bm{\Theta}}})$, has stationary configurations $(\tilde{\bf X}^{*}_{{\bm{\Theta}}},{\bf C}^{*}_{{\bm{\Theta}}})$ satisfying
@@
where $\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}\in\mathbb{R}^{N_{D}\times N\times 3}$ and $\nabla_{{\bm{\Theta}}}{\bf C}^{*}_{{\bm{\Theta}}}\in\mathbb{R}^{N_{D}\times 3\times 3}$ are implicit derivatives that determine how a stationary configuration changes with the variation of potential parameters. In practical applications, variation ${\bf C}$ is typically constrained; for simplicity, we will only consider fixed-volume simulations or isotropic variations controlled by a homogeneous strain $\epsilon^{*}_{{\bm{\Theta}}}\in\mathbb{R}$ around a reference supercell ${\bf C}_{0}$
@@
where $\nabla_{{\bm{\Theta}}}\epsilon^{*}_{{\bm{\Theta}}}\in\mathbb{R}^{N_{D}}$. Taylor expanding (1) to first order in $\delta{{\bm{\Theta}}}$, it is simple to show $\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*},\nabla_{{\bm{\Theta}}}\epsilon_{{\bm{\Theta}}}^{*}$ solve the system of linear equations
@@
where $\nabla^{2}_{\tilde{\bf X}\tilde{\bf X}}{U}$ is the Hessian matrix in scaled coordinates, $\nabla^{2}_{\epsilon\epsilon}{U}$ is proportional to the bulk modulus of the system and $-\nabla^{2}_{\epsilon\tilde{\bf X}}{U}$ is proportional to change in atomic forces under a homogeneous strain and $\nabla^{2}_{{{\bm{\Theta}}}\tilde{\bf X}}{U}$, $\nabla^{2}_{{{\bm{\Theta}}}\epsilon}{U}$ are mixed curvatures. Whilst solution of (5) in principle requires $\mathcal{O}(N^{3})$ effort due to the Hessian, we introduce a novel Hessian-free solution method below, allowing application to large systems.
@@ II.2 Taylor expansion of stationary energies and volumes using implicit derivatives
where ${\bf H}_{\zeta}\in\mathbb{R}^{N_{D}\times N_{D}}$ is a generalized curvature within a given level of approximation and $\zeta={c,h,ih,h+ih}$ is the level of approximation used. Expressions for $\nabla_{{{\bm{\Theta}}}}\epsilon^{*}_{\zeta}$ and ${\bf H}_{\zeta}$ are given in the SM.
Whilst all approaches predict changes in energy, only $\zeta={h,ih,h+ih}$ predict changes in structure. For constant volume relaxations, the inhomogeneous $\zeta=ih$ expansion will be asymptotically exact. For variable volume relaxations, the full $\zeta=h+ih$ expansion will be asymptotically exact, but as we show below the cheaper homogeneous $\zeta=h$ expansion can also give accurate results if we are primarily interested in changes to the energy or volume, rather than structure.
@@ II.3 Evaluation of the implicit derivative through sparse and dense linear algebra methods
In the linear equations (5), the $\nabla^{2}_{\epsilon\epsilon}U$ and $\nabla^{2}_{\epsilon\tilde{\bf X}}U$ derivatives in (5) require only a few $\mathcal{O}(N)$ force calls for evaluation. As a result, the homogeneous approximation requires minimal computational effort, but all knowledge of structural changes is missing as $\nabla_{{\bm{\Theta}}}\tilde{\bf X}^{*}$ is not evaluated. Evaluation of $\nabla_{{\bm{\Theta}}}\tilde{\bf X}^{*}$ for the inhomogeneous approach requires $\mathcal{O}(N^{2})$ finite difference evaluation of the Hessian matrix $\nabla^{2}_{\tilde{\bf X}\tilde{\bf X}}U$ and $\mathcal{O}(N^{3})$ solution of the dense linear equation (5). Whilst of reasonable cost for small systems ($N<2000$), study of extended defects where $10^{4}<N<10^{6}$ requires significant, typically prohibitive, computational resources and careful use of shared memory parallel linear algebra techniques [6].
To overcome this limitation, we note that the Hessian matrix is highly sparse due to the strong locality of atomic forces. In this regime, efficient solutions of the linear equation (5) can be obtained using iterative algorithms such as gmres. In addition, such algorithms do not require access to every element of the Hessian at each iteration, only a linear operator which gives the action of the Hessian on some vector ${\bf V}$, i.e. $\mathcal{L}({\bf V})={\bf V}\nabla^{2}_{\tilde{\bf X}\tilde{\bf X}}U$.
@@
which clearly limits to $\mathcal{L}({\bf V})={\bf V}\nabla^{2}_{\tilde{\bf X}\tilde{\bf X}}{U}$ as desired and only requires $\mathcal{O}(N)$ force calls for evaluation, repeated for each vector $[\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}]_{l}\in\mathbb{R}^{N\times 3}$, $l\in[1,N_{D}]$ of the implicit derivative. Details of our $\mathcal{O}(NN_{D})$ massively parallel method to compute ${\bf\nabla}_{{\bm{\Theta}}}\tilde{\bf X}^{*}$ in LAMMPS [20] through the solution of (5) using (8) at finite values of $\alpha$ is described in methods section IV.1. In section II.8 we apply the method to enable use of the implicit derivative in large atomic systems.
@@ II.4 Evaluation of the implicit derivative with automatic differentiation methods
AD-enabled simulation schemes such as jax-md [16] can clearly evaluate all terms in equation (5) or the sparse operator approach (8). Recently, implicit differentiation schemes have been implemented in jaxopt [17, 18], allowing direct evaluation of e.g. ${\bf\nabla}_{{\bm{\Theta}}}\tilde{\bf X}^{*}$ by differentiating through the minimization algorithm chosen in jax-md. We have implemented and tested all approaches in AD for the binary Lennard-Jones system described below. Despite the simplicity of the potential form, using AD to evaluate terms in (5) or using AD implicit derivative schemes incurs extremely large memory usage, reaching the 80GB limit on A100 GPUs for only a few thousand atoms, as we detail in the supplementary material (SM). We thus conclude that existing AD schemes for direct evaluation of the implicit derivative or Hessian matrices are ill-suited for application to the thousand-atom systems essential for many materials science problems. In contrast, our sparse operator (8) has the same memory usage as any structural minimization. Whilst still incurring a significantly greater memory burden than non-AD methods implemented in LAMMPS, our sparse operator is ideal for implicit derivative evaluation in AD-enabled schemes. Investigation of how (8) can be used with neural network-based interatomic potentials[12] is left for a future study.
Figure 2: Automatic implicit differentiation with classical potentials. a, b Atomic configurations of a single vacancy configuration with $\sigma_{\rm AB}=1$ (unary) and $\sigma_{\rm AB}=1.1$ (random binary). Atom color: fcc centrosymmetry $\in[0,1.25]$. c Relaxed energy for $\sigma_{\rm AB}\in[0.9,1.1]$ for a fixed volume, where constant and homogeneous approximations are identical. The inhomogeneous expansion accounts for changes in $\tilde{\bf X}$, giving much more accurate energy predictions.
Figure 3: Implicit differentiation of vacancy defect formation with SNAP potentials. a, b Formation volume and formation energy vs potential perturbation magnitude for one of 100 tungsten potential samples used in the study. c Average of the absolute value of vacancy formation volume vs perturbation magnitude. On the inset, predicted vacancy formation volume vs its true value obtained with the full relaxation. d Error vs change in vacancy formation energy averaged over the set of potentials and perturbation magnitudes. On the inset, zoom into a smaller formation energy range. In c,d the shaded region indicates the [16%,84%] percentile (±$\pm 1\sigma$) range across all parameter variations. Line colors across the panels indicate the prediction method with the legend presented in a. In d, symbols are added at data points.

### v2

# Exploring parameter dependence of atomic minima with implicit differentiation

@@ Abstract
Interatomic potentials are essential to go beyond ab initio size limitations, but simulation results depend sensitively on potential parameters. Forward propagation of parameter variation is key for uncertainty quantification, whilst backpropagation has found application for emerging inverse problems such as fine-tuning or targeted design. Here, the implicit derivative of functions defined as a fixed point is used to Taylor-expand the energy and structure of atomic minima in potential parameters, evaluating terms via automatic differentiation, dense linear algebra or a sparse operator approach. The latter allows efficient forward and backpropagation through relaxed structures of arbitrarily large systems. The implicit expansion accurately predicts lattice distortion and defect formation energies and volumes with classical and machine-learning potentials, enabling high-dimensional uncertainty propagation without prohibitive overhead. We then show how the implicit derivative can be used to solve challenging inverse problems, minimizing an implicit loss to fine-tune potentials and stabilize solute-induced structural rearrangements at dislocations in tungsten.
Figure 1: Schematic of the implicit derivative approach in atomic systems. a A local minimum ${\bf X}^{*}_{{\bm{\Theta}}}$ will change under changes in potential parameters ${{\bm{\Theta}}}$. The implicit derivative is the gradient of this change, allowing a first-order prediction of structural variations and a second-order prediction of energy variations. b Uncertainty quantification application. Having the uncertainty in the parameter space, $\Delta{{\bm{\Theta}}}$, implicit derivative allows one to compute the uncertainty of any molecular statics property $\mathit{Q}$. c Inverse design application. With initial configuration ${\bf X}^{*}_{{{\bm{\Theta}}}_{1}}$ and potential ${{\bm{\Theta}}}_{1}$, one defines a target configuration ${\bf X}^{*}_{{{\bm{\Theta}}}_{2}}$. Implicit derivative allows one to effectively minimize the potential parameters that provide the target configuration.
@@ I Introduction
Regardless of the interatomic potential employed, changing parameters will change any quantity of interest extracted from a simulation. For simple classical models, parameter variation is essential to explore the model’s phenomenology [7, 5]. For modern data-driven models that target quantitative accuracy, parameter uncertainties can be estimated by Bayesian regression on training data [8, 9] and should be forward-propagated to simulation results to bound quantities of interest [10, 11, 12, 13].
@@
Backpropagation of structural modifications to changes in parameters is finding application in inverse material design [14, 15], training interatomic potentials from experimental data [16], or ‘fine-tuning’ simple interatomic potentials to reproduce desired self-assembly kinetics [7]. More recently, ‘universal’ machine learning potentials have shown near-quantitative accuracy across large portions of the periodic table [17, 18]. This has raised interest in backpropagation for fine-tuning universal models for specific applications [19], and opened the possibility of targeted design through navigation of a smooth latent composition representation [20].
@@
Forward and backpropagation of parameter variation through complex simulations is typically achieved using reverse-mode automatic differentiation (AD) routines, which offer high arithmetic efficiency at the price of large memory requirements  [21, 22, 23, 24]. Whilst AD offers clear advantages in implementation, their significant memory burden complicates application to large atomic systems, especially when targeting higher derivatives beyond forces. Existing methods to propagate variation in parameters to variation in simulation results thus typically employ resampling: new parameters are drawn from some distribution [10] and simulations are repeated. Whilst conceptually straightforward, backpropagation is not possible, assessing convergence is challenging, and the cost can be potentially very large, especially for simulation results requiring geometry minimization for each potential sample.
@@
In this paper, we explore an alternative approach, analytically expanding the structure of relaxed minima to first order in parameter variation, giving a second order expansion of the total energy. The expansion is achieved through evaluation of an implicit derivative [25], i.e. the derivative of a function defined as fixed point, in this case, the atomic structure of a local minimum (Fig. 1a). Our main results are that the implicit derivative enables 1) forward propagation of parameter uncertainties to simulation results for orders of magnitude less computational effort than resampling schemes, allowing rapid propagation of parameter uncertainties and 2) backpropagation of structural variations to target composition-induced structural rearrangements in multi-thousand-atom systems, a challenging task for any other approach (see illustration of these two ideas in Fig. 1b and 1c).
@@
We implement and compare methods to evaluate the implicit derivative using AD [22, 23], dense linear algebra and a sparse operator approach, using the jax-md [21] and LAMMPS [26] simulation codes. We find AD routines for the implicit derivative reach GPU memory limits for $\sim$1000-atom systems even on best-in-class hardware. In contrast, the sparse operator technique reduces to a constrained minimization in LAMMPS, allowing memory-efficient and highly parallelized evaluation. Our developed approach enables uncertainty quantification and inverse design studies with the large atomic simulations essential to capture realistic defect structures.
@@
The paper is structured as follows. We first define the implicit derivative, the Taylor expansion approximations used and their evaluation using automatic differentiation or linear algebra techniques. We then describe how the implicit derivative can be used in forward and backpropagation of parameter variations. Forward propagation of parameter variation is demonstrated using classical and machine learning potentials to explore lattice distortion and vacancy defect formation [34]. Finally, we apply the implicit derivative approach to the inverse material design problems, finding parameter variations that stabilize subtle solute-induced dislocation core reconstructions in tungsten [4].
@@ II.1 Implicit derivative of atomic configurations
We consider a system of $N$ atoms in a periodic supercell of volume $V$, with atomic coordinates ${\bf X}\in\mathbb{R}^{N\times 3}$ and a supercell matrix ${\bf C}\in\mathbb{R}^{3\times 3}$, $V={\rm det}({\bf C})$. Changes to the supercell ${\bf C}\to{\bf C}+\delta{\bf C}$ are defined to induce homogeneous deformations, as in e.g. energy volume curves. We then define scaled atomic coordinates $\tilde{\bf X}$, such that ${\bf X}\equiv\tilde{\bf X}{\bf C}$ and the tuple $(\tilde{\bf X},{\bf C})$ fully describes atomic configurations with periodic boundary conditions. With a vector of $N_{D}$ potential parameters ${{\bm{\Theta}}}$, a potential energy model ${U}(\tilde{\bf X},{\bf C};{{\bm{\Theta}}})$, has stationary configurations $(\tilde{\bf X}^{*}_{{\bm{\Theta}}},{\bf C}^{*}_{{\bm{\Theta}}})$ satisfying
@@
where $\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}\in\mathbb{R}^{N_{D}\times N\times 3}$ and $\nabla_{{\bm{\Theta}}}{\bf C}^{*}_{{\bm{\Theta}}}\in\mathbb{R}^{N_{D}\times 3\times 3}$ are implicit derivatives that determine how a stationary configuration changes with the variation of potential parameters. In practical applications, supercell variation ${\bf C}$ is typically constrained; for simplicity, we will only consider fixed-volume simulations or isotropic variations controlled by a homogeneous strain $\epsilon^{*}_{{\bm{\Theta}}}\in\mathbb{R}$ around a reference supercell ${\bf C}_{0}$
@@
where $\nabla_{{\bm{\Theta}}}\epsilon^{*}_{{\bm{\Theta}}}\in\mathbb{R}^{N_{D}}$. Taylor-expanding equations (1) to first order in $\delta{{\bm{\Theta}}}$, it is simple to show that $\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*},\nabla_{{\bm{\Theta}}}\epsilon_{{\bm{\Theta}}}^{*}$ solve the system of linear equations
@@
where $\nabla^{2}_{\tilde{\bf X}\tilde{\bf X}}{U}$ is the Hessian matrix in scaled coordinates, $\nabla^{2}_{\epsilon\epsilon}{U}$ is proportional to the bulk modulus of the system, $-\nabla^{2}_{\epsilon\tilde{\bf X}}{U}$ is proportional to change in atomic forces under a homogeneous strain and $\nabla^{2}_{{{\bm{\Theta}}}\tilde{\bf X}}{U}$, $\nabla^{2}_{{{\bm{\Theta}}}\epsilon}{U}$ are mixed curvatures. Whilst the solution of (5) in principle requires $\mathcal{O}(N^{3})$ effort due to the Hessian, we introduce a Hessian-free solution method below, allowing application to large systems.
@@ II.2 Taylor expansion of stationary energies and volumes using implicit derivatives
where ${\bf H}_{\zeta}\in\mathbb{R}^{N_{D}\times N_{D}}$ is a generalized curvature within a given level of approximation and $\zeta={c,h,ih,h+ih}$ is the level of approximation used. Expressions for $\nabla_{{{\bm{\Theta}}}}\epsilon^{*}_{\zeta}$ and ${\bf H}_{\zeta}$ are given in the supplementary material.
In all three approaches, the potential parameters ${{\bm{\Theta}}}$ vary; therefore, each method predicts changes in energy. However, only $\zeta={h,ih,h+ih}$ predict changes in structure. For constant volume relaxations, the inhomogeneous $\zeta=ih$ expansion is asymptotically exact. For variable volume relaxations, the full $\zeta=h+ih$ expansion is asymptotically exact, but as we show below, the cheaper homogeneous $\zeta=h$ expansion can also give accurate results when the focus is on changes to the energy or volume, rather than structure.
@@ II.3 Evaluation of the implicit derivative through sparse and dense linear algebra methods
In the linear equations (5), the $\nabla^{2}_{\epsilon\epsilon}U$ and $\nabla^{2}_{\epsilon\tilde{\bf X}}U$ derivatives require only a few $\mathcal{O}(N)$ force calls for evaluation. As a result, the homogeneous approximation requires minimal computational effort, but all knowledge of structural changes is missing as $\nabla_{{\bm{\Theta}}}\tilde{\bf X}^{*}$ is not evaluated. Evaluation of $\nabla_{{\bm{\Theta}}}\tilde{\bf X}^{*}$ for the inhomogeneous approach requires $\mathcal{O}(N^{2})$ finite difference evaluations of the Hessian matrix $\nabla^{2}_{\tilde{\bf X}\tilde{\bf X}}U$ and $\mathcal{O}(N^{3})$ solution of the dense linear equation (5). Whilst of reasonable cost for small systems ($N<2000$), study of extended defects where $10^{4}<N<10^{6}$ requires significant, typically prohibitive, computational resources and careful use of shared memory parallel linear algebra techniques [6].
To overcome this limitation, we note that the Hessian matrix is highly sparse due to the strong locality of atomic forces. In this regime, efficient solutions of the linear equations (5) can be obtained using iterative algorithms. In addition, such algorithms do not require access to every element of the Hessian at each iteration, only a linear operator that gives the action of the Hessian on some vector ${\bf V}\in\mathbb{R}^{N\times 3}$, i.e. $\mathcal{L}({\bf V})={\bf V}\nabla^{2}_{\tilde{\bf X}\tilde{\bf X}}U$.
@@
which, in the limit, is equal to $\mathcal{L}({\bf V})={\bf V}\nabla^{2}_{\tilde{\bf X}\tilde{\bf X}}{U}$ as desired and only requires $\mathcal{O}(N)$ force calls for evaluation. To compute the inhomogeneous implicit derivative, we apply the sparse linear operator to each vector $[\nabla_{{\bm{\Theta}}}\tilde{\bf X}_{{\bm{\Theta}}}^{*}]_{l}\in\mathbb{R}^{N\times 3}$, $l\in[1,N_{D}]$. Details of our $\mathcal{O}(NN_{D})$ massively parallel method to compute ${\bf\nabla}_{{\bm{\Theta}}}\tilde{\bf X}^{*}$ in LAMMPS [26] at finite values of $\alpha$ is described in methods section IV.1. In section II.8, we apply the method to enable the use of the implicit derivative in large atomic systems.
@@ II.4 Evaluation of the implicit derivative with automatic differentiation methods
AD-enabled simulation schemes such as jax-md [21] can clearly evaluate all terms in equation (5) or the sparse operator approach (8). Recently, implicit differentiation schemes have been implemented in jaxopt [22, 23], allowing direct evaluation of e.g., ${\bf\nabla}_{{\bm{\Theta}}}\tilde{\bf X}^{*}$ by differentiating through the minimization algorithm chosen in jax-md. We have implemented and tested all approaches in AD for the binary Lennard-Jones system described below. Despite the simplicity of the potential form, using AD implicit derivative schemes incurs extremely large memory usage, reaching the 80GB limit on NVIDIA A100 GPUs for only a few thousand atoms, as we detail in the supplementary material. We thus conclude that existing AD schemes for direct evaluation of the implicit derivative or Hessian matrices are ill-suited for application to the thousand-atom systems essential for many materials science problems. In contrast, our sparse operator technique has the same memory usage as any structural minimization. Whilst still incurring a significantly greater memory burden than non-AD methods implemented in LAMMPS, our sparse operator is ideal for implicit derivative evaluation in AD-enabled schemes.
Figure 2: Automatic implicit differentiation with classical potentials. a, b Atomic configurations of a single vacancy configuration with $\sigma_{\rm AB}=1$ (unary) and $\sigma_{\rm AB}=1.1$ (random binary). Atom colors encode fcc centrosymmetry in the range $[0,1.25]$; black color corresponds to perfect fcc structure. c Relaxed energy and lattice distortion for $\sigma_{\rm AB}\in[0.9,1.1]$ at fixed volume. The true values are obtained with the energy minimization at each $\sigma_{\rm AB}$ value.
Figure 3: Implicit differentiation of vacancy defect formation with SNAP potentials. a, b Formation volume and formation energy vs potential perturbation magnitude for one of 100 tungsten potential samples used in the study. c Average absolute error of formation volume vs perturbation magnitude. On the inset, predicted vacancy formation volume vs its true value obtained with the full relaxation. d Error vs change in vacancy formation energy averaged over the set of potentials and perturbation magnitudes. On the inset, zoom into a smaller formation energy range. In c,d the shaded regions indicates the [16%, 84%] percentile (±$\pm 1\sigma$) range across all parameter variations. Line colors across the panels indicate the prediction method with the legend presented in a. In d, symbols are added at data points. Pristine supercell size of 128 atoms was used in the study.